OpenStudy (cria7411):

Trevor solved the system of equations below. What mistake did he make in his work? 2x + y = 5 x − 2y = 10 y = 5 − 2x x − 2(5 − 2x) = 10 x − 10 + 4x = 10 5x − 10 = 10 5x = 0 x = 0 2(0) + y = 5 y = 5 A) He should have substituted 5 + 2x B) He combined like terms incorrectly, it should have been 4x instead of 5x C) He subtracted 10 from the right side instead of adding 10 to the right side D) He made no mistake

OpenStudy (tehhota):

JESUS BUS IS REAL AND ITS BACK

OpenStudy (mecharv):

Can you please try to type the equations clearly again? they are super confusing

OpenStudy (da-jesus-train):

Da-Jesus-Train is real and Back

OpenStudy (tehhota):

JESUS Planes Trains and Automobiles

OpenStudy (tehhota):

JESUS JESUS JESUS

OpenStudy (cria7411):

Trevor solved the system of equations below. What mistake did he make in his work? 2x + y = 5 x − 2y = 10 y = 5 − 2x x − 2(5 − 2x) = 10 x − 10 + 4x = 10 5x − 10 = 10 5x = 0 x = 0 2(0) + y = 5 y = 5 A) He should have substituted 5 + 2x B) He combined like terms incorrectly, it should have been 4x instead of 5x C) He subtracted 10 from the right side instead of adding 10 to the right side D) He made no mistake

OpenStudy (mecharv):

Yes now it is very clear and I found out the mistake... where do you think he made a mistake?

OpenStudy (cria7411):

B

OpenStudy (mecharv):

Nop now you made a mistake

OpenStudy (cria7411):

Is it A?

OpenStudy (tehhota):

ITS JESUS

OpenStudy (tehhota):

ITS THE JESUS BUS

OpenStudy (tehhota):

JESUS CARES ABOUT MATH

OpenStudy (cria7411):

Hello you still there @mecharv

OpenStudy (mecharv):

Yes it is not A either... see 5x-10=10 then, 5x=20 right??

OpenStudy (cria7411):

So it is C right?

OpenStudy (mecharv):

Perfect :) But make sure you get better hold of the concept :)

OpenStudy (morezuela):

Yes, but do you see why?

OpenStudy (cria7411):

Of course.

OpenStudy (morezuela):

It's a good practice to explain now why C is correct and why A, B and C are not. This way you'll get a good understanding of the concepts involved. Try it.